# A Stochastic Approach to the Definition of the Path Integral Measure

### v2

@@ 5.2 Semigroup structure and Riemann Product
In the case of the Feyman-Kac theorem and all of the existing rigorous formulations of the Path Integral, the formulae define Markov semigroups via their diffusion generators. In the bundle framework, the bundle law captures all the marginas defined for the same $t\in[0,1]$, and hence to represent the amplitudes over a multitude of points, we may define a transition kernel $K(Q_{x},Q_{y})$ between two configurations, which given the stochasticity of the diffusion process in $\mathcal{T}_{\eta}$ is a Markov kernel. Alongside the kernel, we may also define a time-dependent generator of $X$ in the parallel frame of the coordinates for the tube with $G_{t}f=\langle b_{t},\nabla f\rangle+\frac{\sigma^{2}}{2}\nabla^{2}f$.
@@ 6 Equivalence of the Stochastic Path Integral and Euclidean Theory
The present section will be dedicated to establishing and proving equivalence between the stochastic framework and the existing rigorous formulations of the path integral realized under Wick’s theorem, which is valid for the Stochastic Path Integral as well. During the writing of this paper, the notation for the tubular bundle changed from $\mathcal{T}_{\eta}$ to $T_{\eta}$ in an attempt to simplify the typing process. We make this notation change explicit and the author apologizes for any and all inconvenience caused by this.
@@ 6.2 Stochastic Path Integral and the Feynman-Kac theorem
Since $f$ is bounded, $\exists M_{f}<\infty$ with $|f(X_{T})|\leq M_{f}$. For $\theta=-i$ we obtain that $\Big|f(X_{T})e^{iA}\Big|=|f(X_{T})|\,\big|e^{iA}\big|=|f(X_{T})|\leq M_{f},$ since $|e^{iA}|=1\quad\forall A\in\mathbb{R}$. In particular, the integrand in for the expectation under $\theta=-i$ is bounded and hence integrable with respect to some probability measure $\mathbb{P}_{0,x}$. Therefore the expectation exists and is finite, and the stochastic Feynman weight $\exp({i\int_{0}^{T}V_{E}})$ Lebesgue integral without necessitating an oscillatory regulation. Assuming that $V_{E}=L_{E}$ where $L_{E}$ is the Euclidean Lagrangian density. Under Wick rotation (as fixed in the main text), the Euclidean action $\int_{0}^{T}L_{E}\,dt$ corresponds to the action $S(X)=\int_{0}^{T}L(t,X_{t},\dot{X}_{t})\,dt$, and we have
@@ 7 Conclusion and Prospective directions
We have successfully formulated a restricted path space that encompasses all physically meaningful trajectories which may be endowed with a Hilbert space bundle structure and a stochastic process, yielding a path integral that does not face the issues of divergence plaguing the heuristic path integral with an undefined measure $\mathcal{D}x(t)$. The framework as detailed in the current paper, however, is still very raw and bound to the preliminary conditions detailed in the introduction.

### v3

@@ 4.2 Semigroup structure and Riemann Product
In the case of the Feyman-Kac theorem and all of the existing rigorous formulations of the Path Integral, the formulae define Markov semigroups via their diffusion generators. In the bundle framework, the bundle law captures all the marginals defined for the same $t\in[0,1]$, and hence to represent the amplitudes over a multitude of points, we may define a transition kernel $K(Q_{x},Q_{y})$ between two configurations, which given the stochasticity of the diffusion process in $\mathcal{T}_{\eta}$ is a Markov kernel. Alongside the kernel, we may also define a time-dependent generator of $X$ in the parallel frame of the coordinates for the tube with $G_{t}f=\langle b_{t},\nabla f\rangle+\frac{\sigma^{2}}{2}\nabla^{2}f$.
@@ 5 Equivalence of the Stochastic Path Integral and Euclidean Theory
The present section will be dedicated to establishing and proving equivalence between the stochastic framework and the existing rigorous formulations of the path integral realized under Wick’s theorem, which is valid for the Stochastic Path Integral as well. During the writing of this paper, the notation for the tubular bundle changed from $\mathcal{T}_{\eta}$ to $T_{\eta}$ in an attempt to simplify the typing process. We make this notation change explicit and the author apologizes for any and all inconvenience caused by this. Since the Feynman-Kac Path Integral also is based on stochastic dynamics, in some cases the path integral from 4.5 will be referred to as the ’tube-confined’.
@@ 5.2 Stochastic Path Integral and the Feynman-Kac theorem
Since $f$ is bounded, $\exists M_{f}<\infty$ with $|f(X_{T})|\leq M_{f}$. For $\theta=-i$ we obtain that $\Big|f(X_{T})e^{iA}\Big|=|f(X_{T})|\,\big|e^{iA}\big|=|f(X_{T})|\leq M_{f},$ since $|e^{iA}|=1\quad\forall A\in\mathbb{R}$. In particular, the integrand in for the expectation under $\theta=-i$ is bounded and hence integrable with respect to some probability measure $\mathbb{P}_{0,x}$, wherefrom the expectation exists and is finite, while the stochastic Feynman weight $\exp({i\int_{0}^{T}V_{E}})$ in the integrand admits a finite-dimensional Lebesgue integral without necessitating an oscillatory regulation. Assuming that $V_{E}=L_{E}$ with $L_{E}$ the Euclidean Lagrangian density, we get that under Wick rotation (as fixed in the main text), the Euclidean action $\int_{0}^{T}L_{E}\,dt$ corresponds to the action $S(X)=\int_{0}^{T}L(t,X_{t},\dot{X}_{t})\,dt$, and we obtain
@@
6 Derivation of the Free Particle Propagator
From a theoretical standpoint, the framework constructed above provides a robust measure and an object for path space compactification for a free particle with a quadratic Lagrangian traveling between two points with finite energy uncertainty and time of travel. It is not necessary to verify whether the framework holds by computing the propagator for a free particle with mass $m$ and a base-trajectory $\gamma_{0}(t)=A+\frac{t}{T}(B-A):T<\infty,t\in[0,1]$ in $\mathbb{R}^{n}$.
The trajectories in the tube are given by $X_{t}=\gamma_{0}(t)+Y_{t}:Y_{0}=Y_{T}=0$, with the action for the system being defined by the integral
$S(\gamma)=\frac{m}{2}\int_{0}^{T}\dot{\gamma}(t)^{2}dt$ (6.1)
From $\dot{\gamma}_{0}(t)=\frac{|B-A|}{T}(1-t),$ we find that $S(\gamma_{0})=\frac{m}{2T}|B-A|^{2}\implies S(\gamma)=\frac{m}{2T}|B-A|^{2}+\frac{m}{2T}\int_{0}^{T}|\dot{Y}_{t}|^{2}dt$. From this, the propagator for the function is simply
$K(B,T;A,0)=\exp(\frac{im}{2T\hbar}|B-A|^{2})\times\mathbb{E}_{\mu}[\exp(\frac{im}{2\hbar}\int_{0}^{T}|\dot{Y}_{t}|^{2}dt)]$ (6.2)
We have previously seen that by Girsanov, we may define a Gaussian measure $\Gamma$ on $H^{1}_{0}([0,T];\mathbb{R}^{n})$ satisfying (4.1). Then, under $\Gamma$, $Y\in H^{1}_{0}:\int_{0}^{T}|\dot{Y}_{t}|^{2}dt<\infty\implies\exp(\frac{im}{2\hbar}\int_{0}^{T}|\dot{Y}_{t}|^{2}dt)=F(Y):|F(Y)|=1,\quad F\in L^{1}(\Gamma)\cap L^{\infty}(\Gamma).$ We may remove the prefactor of $\frac{im}{2\hbar}$ and just view this expression through the lens of a random variable under $\Gamma$,
$Q(Y):=\int_{0}^{T}|\dot{Y}_{t}|^{2}\,dt$ (6.3)
This is a quadratic form on the real Hilbert space $H:=H^{1}_{0}([0,T];\mathbb{R}^{n})$ with a corresponding inner product $\langle Y,Z\rangle_{H}:=\int_{0}^{T}\dot{Y}_{t}\cdot\dot{Z}_{t}\,dt.$ Meanwhile, the Gaussian measure $\Gamma$ has covariance operator $C=(-\Delta_{D})^{-1},$ where $\Delta_{D}$ is the Dirichlet Laplacian on $[0,T]$. Then, let there be some family of functions $\{e_{k}(t)\}$ with
$e_{k}(t)=\sqrt{\frac{2}{T}}\sin\!\left(\frac{k\pi t}{T}\right),\qquad\lambda_{k}=\left(\frac{k\pi}{T}\right)^{2}.$ (6.4)
Then, under $\Gamma$,
$Y(t)=\sum_{k=1}^{\infty}\xi_{k}e_{k}(t),$ (6.5)
with $\xi_{k}\sim\mathcal{N}\!\left(0,\lambda_{k}^{-1}\right)$ defined component-wise in $\mathbb{R}^{n}$. We may then compute $Q(Y)=\sum_{k=1}^{\infty}\lambda_{k}\,|\xi_{k}|^{2},$ obtaining that
$\mathbb{E}_{\Gamma}\!\left[e^{\frac{im}{2\hbar}Q(Y)}\right]=\prod_{k=1}^{\infty}\mathbb{E}\!\left[\exp\!\left(\frac{im}{2\hbar}\lambda_{k}|\xi_{k}|^{2}\right)\right]^{n}.$ (6.6)
For a Gaussian random variable over $\mathbb{R}$, $\xi\sim\mathcal{N}(0,\lambda^{-1})$, $\mathbb{E}\!\left[e^{ia\lambda\xi^{2}}\right]=(1-2ia)^{-1/2},\qquad a\in\mathbb{R}.$ Hence,
$\mathbb{E}\!\left[e^{\frac{im}{2\hbar}\lambda_{k}\xi_{k}^{2}}\right]=\left(1-\frac{im}{\hbar}\right)^{-1/2}.$ (6.7)
Taking the infinite product yields
$\mathbb{E}_{\Gamma}\!\left[e^{\frac{im}{2\hbar}Q(Y)}\right]=\prod_{k=1}^{\infty}\left(1-\frac{im}{\hbar}\right)^{-n/2}.$ (6.8)
This product is understood in the zeta–regularized sense, giving the determinant representation
$\mathbb{E}_{\Gamma}\!\left[e^{\frac{im}{2\hbar}Q(Y)}\right]=\det\nolimits^{-n/2}\!\left(I-\frac{im}{\hbar}(-\Delta_{D})^{-1}\right).$ (6.9)
By leveraging the classical identity $\prod_{k=1}^{\infty}\left(1+\frac{a^{2}}{k^{2}\pi^{2}}\right)=\frac{\sinh a}{a},$ together with analytic continuation $a\mapsto e^{-i\pi/2}a$, we obtain
$\mathbb{E}_{\Gamma}\!\left[\exp\!\left(\frac{im}{2\hbar}\int_{0}^{T}|\dot{Y}_{t}|^{2}\,dt\right)\right]=\left(\frac{m}{2\pi i\hbar T}\right)^{\frac{n}{2}}$ (6.10)
This last expression is a genuine $L^{1}(\Gamma)$ expectation, not a distributional limit, and it is evaluated entirely in real (Lorentzian) time. Therefore, under Gaussian time, we find a result equal to the results of the traditional path integral for the free particle.
@@ 7 Conclusion and Prospective directions
We have formulated a restricted path space that encompasses all physically meaningful trajectories which may be endowed with a Hilbert space bundle structure and a stochastic process, yielding a path integral that does not face the issues of divergence plaguing the heuristic path integral with an undefined measure $\mathcal{D}x(t)$. The framework as detailed in the current paper, however, is still very raw and bound to the preliminary conditions detailed in the introduction.
@@
Appendix A Appendix: On Relevant Trajectories
$f(E)=\frac{1}{(\delta E)^{2}-(\Delta E)^{2}}\implies f(H)=\frac{1}{(\delta E)^{2}-(H-E_{0})^{2}}$ (A.1)
$\Phi(\gamma,\gamma_{0})=\oint_{\gamma\circ\gamma_{0}}f(H)p_{i}dq^{i}$ (A.2)
For any $\gamma:|H(\gamma)-E_{0}|>\delta E,$ the contour approaches a pole of $f(H)$, wherefrom by the Residue Theorem $\Phi(\gamma,\gamma_{0})$ diverges. The class of admissible trajectories is therefore given by the homotopies of $\gamma_{0}$ for which $\Phi$ is finite. This function is not rigorous, but is useful for heuristic intuition motivating the restriction of the path space.